OpenStudy (anonymous):

winner of a 5 mile bike race finishes 5 minutes ahead of a teammate and travels 5mph

OpenStudy (anonymous):

average speed of each racer?

OpenStudy (anonymous):

winner takes 5mph/5mile = 1hr to complete the race thus the loser takes 1hr+(5/60)hr =1+1/12=13/12 hr to complete the 5 mile race thus the average speed of the loser is 5/(13/12) =60/13 mph=appax 4.62 mph